Here's my question: Bob invested $20,000, part at 14% and part at 13%. If the total interest at the end of the year is $2,720, how much did he invest at 14%? I got $11,000, but I don't think that's right.
Answer by checkley71(8403)

.14X+.13(20000-X)=2720
.14X+2600-.13X=2720
.14X-.13X=2720-2600
.01X=120
X=120/.01
X=$12,000 AMOUNT INVESTED @ 14%.
20000-12000=8000 INVESTED @ 13%
PROOF
.14*12000+.13*8000=2720
1680+1040=2720
2720=2720